OpenStudy (anonymous):

Need help with this real world problem. "A 65 foot rope hangs from the top of a pole. When pulled taut to it's fullest length the rope reaches a point on the ground 30 feet from the base of the pole. Find the height of the pole to the NEAREST TENTH of a foot. Let P equal the height of the pole. "Use the Pythagorean Theorem". So far I have: a^2+b^2=c^2 30^2+p^2=65^2 900+p^2=4,225 p^2=4,225-900 p^2=3,325 p= ?????? I can't seem to get that last part about finding the height of the pole to the nearest tenth of a foot.

OpenStudy (anonymous):

Do you know what is 'tenth'?

OpenStudy (anonymous):

No I don't. I just know that it needs to be put into a decimal form. I think? Not really sure at all. :\

OpenStudy (anonymous):

Tenth is the first digit to the right of the decimal point.

OpenStudy (anonymous):

57.6 (Truncating)

OpenStudy (anonymous):

But how am I supposed to figure that out with 3,325? That is what confuses me. I can't figure out what P is unless I know how to turn 3,325 into a decimal. If that makes any sense?

OpenStudy (anonymous):

just square root ur 3325

OpenStudy (anonymous):

give u decimal answer

OpenStudy (anonymous):

That would give me 11055625?

OpenStudy (anonymous):



OpenStudy (anonymous):

Like for example. I have this one p^2 = 3844 - 961 p^2 = 2,883 p = 53.226 = 53.2 How did they come up with 53.226? I'm not that great with the whole square root thing.

OpenStudy (anonymous):

calculator ur best pal in this situation